# Networks of neural networks: more is different

### v1

@@ 3 Disentangling spurious states
The modular structure of an $L$-directional associative memory, can be leveraged to tackle several kinds of task beyond standard pattern retrieval. For instance, in AABCR-2024 , we considered pattern disentanglement in the case where patterns retrievable by different layers were independent. Dropping this independence condition makes the task more challenging and, in the current work, we deepen such a scenario. Specifically, we aim to exploit the neural network introduced in the previous section for disentangling spurious states, that is, we want to input information in the form of a mixtures of $L$ patterns (without loss of generality we consider the first $L$ patterns) as $\textrm{sign}(\bm{\xi}^{1}+\bm{\xi}^{2}+...+\bm{\xi}^{L})$ and to get as output the single components $\bm{\xi}^{1},\bm{\xi}^{2},...,\bm{\xi}^{L}$, one per layer. In other words, we want the configurations $\bm{\sigma}^{a}=\bm{\xi}^{a}$ for $a=1,...,L$ (or any permutation that ensures that different layers retrieve all the different patterns in the input), referred to as $\bm{\sigma}^{(1,2,...,L)}$, to be stable and attracting the configuration $\bm{\sigma}^{a}=\textrm{sgn}(\bm{\xi}^{1}+\bm{\xi}^{2}+...+\bm{\xi}^{L})$ for $a=1,...,L$. Given this task, a natural choice for the field acting on each layer is
$h_{i}=\mathrm{sign}\left(\sum_{\mu=1}^{L}\xi_{i}^{\mu}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ \textrm{for}\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,...,N\leavevmode\nobreak\ \leavevmode\nobreak\ \textrm{and}\leavevmode\nobreak\ \leavevmode\nobreak\ a=1,...,L$ (12)
as this is the only available information for the user; notice that this field is layer independent.
The evolution towards the target configuration $\bm{\sigma}^{(1,2,...,L)}$ can be checked by different means. In particular, in Secs. 3.1-3.2, we analytically investigate whether the latter corresponds to the equilibrium configuration resulting from the self-consistent equations (11), when the fields (12) are applied. Next, in Secs. 3.3-3.4, we numerically investigate whether, starting from the input configuration $\bm{\sigma}^{a}=\bm{h}$ for $a=1,...,L$ and applying the stochastic local-field-alignment (see, e.g., Amit ), the system eventually reaches the target configuration and this is stable. We recall that the stochastic local-field-alignment plays as neural dynamics for the network and reads
@@
$\tilde{h}_{i}^{a}(t)$ $=$ $\dfrac{1}{N}\displaystyle\sum\limits_{b=1}^{L}g_{ab}\displaystyle\sum\limits_{\mu=1}^{K}\displaystyle\sum\limits_{j=1}^{N}\xi_{j}^{\mu}\sigma_{j}^{b}(t)\xi_{i}^{\mu}+Hh^{a}_{i}$ (14)
where $t$ denotes the time step, $\zeta_{i}^{a}(t)$ is a stochastic contribution [^5] We will set $\zeta=\textrm{atanh}(x)$ with $x$ a uniform random variable ranging in $[-1,+1]$; this choice ensures that the dynamics (13) yields to a Boltzmann-Gibbs equilibrium, such that this network can be seen as a generalized Boltzmann machine Coolen . and $\bm{\tilde{h}}^{a}\in\mathbb{R}^{N}$ is the local field acting on neurons in the $a$-th layer (stemming from the interactions with other neurons and from the external field).
@@ 3.1 Stability analysis in the high-load, noiseless regime
As mentioned in Sec. 2, the configuration where different layers retrieve different patters is only one (out of many) possible extrema for the Hamiltonian (3). Thus, before inspecting the ability of the model to disentangle spurious states, it is worth taking a look at some representative extremal configurations and at their stability in the noiseless scenario ($\beta\to\infty$). We set $L=3$ (we refer to App. C for an analysis of the case $L=5$ which confirms the result robustness) and we focus on the following classes of neuronal configurations [^6] We are referring to “classes” of neural configurations, because, beyond the degeneracy due to the permutation of the three patterns over the three layers, there is also a degeneracy due to the symmetry of the Hamiltonian (5) under spin flip of all the three layers.:
@@
For each of them we will estimate the energy, the consistency and the stability. Before proceeding, a couple of remarks are in order. First, the previous neuronal states have been chosen because they are recognized to minimize at least one of the contributions making up the Hamiltonian (3) and, in fact, we checked that they are also solutions of the self-consistency equations. However, we recall that the last condition only ensures that these configurations are extremal for the free energy, but not necessarily minima, that is, stable points: this prompts the study of the free-energy Hessian performed in Sec. 3.2. Second, here the consistency analysis is pursued by recalling the stochastic dynamics (13) and checking whether the configurations remain unchanged, that is, recasting (13) into an evolutionary rule for the Mattis magnetizations
$m^{a}_{\mu}(t+1)=\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\xi_{i}^{\mu}\sigma_{i}^{a}(t+1)=\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\xi_{i}^{\mu}\sigma_{i}^{a}(t)\mathrm{sign}\left[\tilde{h}_{i}^{a}\sigma_{i}^{a}(t)\right]\leavevmode\nobreak\ \leavevmode\nobreak\ \textrm{for}\leavevmode\nobreak\ \leavevmode\nobreak\ a=1,...,L,$ (15)
@@
Let us start with the analytical inspection:
  • $\bm{\sigma}^{(1,2,3)}=(\bm{\xi}^{1},\bm{\xi}^{2},\bm{\xi}^{3})$
    This is our target configuration, whose magnetization is $m_{\mu}^{a}=\delta_{\mu a}$ for $a=1,...,3$ (apart from vanishing corrections in the thermodynamic limit). This configuration minimizes the first contribution in the Hamiltonian (5), whose value can be estimated in the large size limit (we exploit the Rademacher nature of pattern entries and the central limit theorem, c.l.t.) to get
    $\frac{\mathcal{H}(\bm{\sigma}^{(1,2,3)})}{N}\underset{c.l.t.}{\sim}-3(1+\gamma)-\frac{3}{2}H+x\frac{\mathcal{C}^{(1,2,3)}}{\sqrt{N}},$ (16)
    $m_{1}^{1}(t=1)=m_{2}^{2}(t=1)=m_{3}^{3}(t=1)=\mathrm{erf}\left[\dfrac{2+H}{\sqrt{2\left(4\gamma+8\lambda^{2}(1+\gamma)-8\lambda H+3H^{2}\right)}}\right].$ (17)
@@
    Figure 2: We initialize the system in a configuration obtained from $\bm{\sigma}^{(1,2,3)}$ (first line), from $\bm{\sigma}^{(1,1,1)}$ (second line), from $\bm{\sigma}^{(1,1,1^{\prime})}$ (third line), and from $\bm{\sigma}^{(h)}$ (fourth line), by flipping randomly its entries – the flip is implemented by multiplying each neuron variable $\sigma_{i}^{a}$ by a random variable $\chi_{i}^{a}$ drawn from $P(\chi)=\frac{1+r}{2}\delta(\chi-1)+\frac{1-r}{2}\delta(\chi+1)$, where $r=1.0$ (left column), $r=0.8$ (middle column), and $r=0.5$ (right column), clearly, the larger $r$ and the closer the initial configuration to the reference. Then, we implement the dynamics (13) with $T=0$, up to convergence to a fixed point. This is repeated for several choices of the parameters $H$ and $\lambda$ sampled uniformly in, respectively, $[0,2]$ and $[0,0.5]$ and for fixed $N=5000$ and $K=50$. Different final states are recorded and represented by different symbols and colors, as reported by the legend on the right: $\bm{\sigma}^{(1,2,3)}$ (green $\times$), $\bm{\sigma}^{(1,1,1)}$ (blue $+$), $\bm{\sigma}^{(1,1,1^{\prime})}$ (magenta $\circ$), $\bm{\sigma}^{(h)}$ (red $\triangle$), or none of those considered in this section (gray $\bullet$). The stability range for the four examined configurations, predicted analytically by studying the Mattis magnetization evolution (17), is represented by the black lines. Specifically, these are obtained by determining in which region of the ($\lambda$, $H$) plane the error functions in, respectively, eqs. (17), (19), (21), (22), (24), exceed a certain threshold, which we set to $0.95$; for $\bm{\sigma}^{(h)}$ no boundaries are detected in the region under consideration. The shade in the color accounts for the energy associated to the related fixed point: the smaller the energy and the darker the color. Thus, for small $H$, although $\bm{\sigma}^{(1,2,3)}$ turns out to be stable, its energy is relatively close to zero.
    As expected, in the limit $H,\lambda\to 0$, (17) recovers the expression for the next-step magnetization of three independent Hopfield models, each initialized with the respective initial condition $\bm{\sigma}^{(a)}=\bm{\xi}^{a}$, $a=1,2,3$.
  • $\bm{\sigma}^{(1,1,1)}\equiv(\bm{\xi}^{1},\bm{\xi}^{1},\bm{\xi}^{1})$
    This state corresponds to the pure retrieval in a standard Hopfield model and minimizes the first contribution in the Hamiltonian (5). Its intensive energy in the large-size limit is
    $\frac{\mathcal{H}(\bm{\sigma}^{(1,1,1)})}{N}\underset{c.l.t.}{\sim}-3(1-\lambda)(1+\gamma)-\frac{3}{2}H+x\frac{\mathcal{C}^{(1,1,1)}}{\sqrt{N}}.$ (18)
    As expected, when $\lambda$ is increased, this configuration makes the coupling between layers more and more frustrated, consequently, the energy grows and the stability is impaired; if $\lambda=0$ the above energy recovers the previous one for $\bm{\sigma}^{(1,2,3)}$. The next-step magnetization in the thermodynamic limit is
    $m_{1}^{1}(t=1)=m_{2}^{2}(t=1)=m_{3}^{3}(t=1)=\mathrm{erf}\left[\dfrac{2(1-2\lambda)+H}{\sqrt{2\left(4\gamma(1-2\lambda)^{2}+3H^{2}\right)}}\right].$ (19)
  • $\bm{\sigma}^{(1,1,1^{\prime})}\equiv(\bm{\xi}^{1},\bm{\xi}^{1},-\bm{\xi}^{1})$
    This staggered configuration minimizes both the first and the third contribution of the Hamiltonian (5). The intensive energy is
    $\frac{\mathcal{H}(\bm{\sigma}^{(1,1,1^{\prime})})}{N}\underset{c.l.t.}{\sim}-(3+\lambda)(1+\gamma)-\dfrac{H}{2}+x\frac{\mathcal{C}^{(1,1,1^{\prime})}}{\sqrt{N}}.$ (20)
    By comparing this expression with (16), (18) and the following (23), we see that, when $H=0$ and $\lambda\neq 0$, this state is the one with the lowest energy among those considered here, in fact, this configuration favors all the intra-layer interactions and partially favours inter-layer interactions. However, by comparing this energy with the one obtained for $\bm{\sigma}^{(1,2,3)}$, we see that there exists a range of values for the parameters $H\neq 0$ and $\lambda$, such that the energy of this state is larger and therefore energetically less convenient.
    $m_{1}^{1}(t=1)=m_{1}^{2}(t=1)=\mathrm{erf}\left[\dfrac{2+H}{\sqrt{2\left(4\gamma+3H^{2}\right)}}\right],$ (21)
    while for the third layer
    $m_{1}^{3}(t=1)=-\mathrm{erf}\left[\dfrac{2+4\lambda-H}{\sqrt{2\left[4\gamma(1+2\lambda)^{2}+3H^{2}\right]}}\right].$ (22)
    Notice that, if $H=0$ and $\gamma\ll 1$, $m_{1}^{1}(t=1)=m_{1}^{2}(t=1)\approx 1$ and their expression recovers the one of a pure state in a standard Hopfield model. Further, if $\lambda\neq 0$, $|m_{1}^{3}(t=1)|$ is as well close to 1 and it is enhanced by $\lambda$ (in fact, the denominator is always smaller than $1$ if $20<\lambda<1/$ 2).
  • $\bm{\sigma}^{(h)}\equiv(\bm{h},\bm{h},\bm{h})$
    This state corresponds to the input mixture, repeated over all the layers. In the large $N$ limit the related intensive energy is
    $\frac{\mathcal{H}(\bm{\sigma}^{(h)})}{N}\underset{c.l.t.}{\sim}-3(1-\lambda)\left(\dfrac{3}{4}+\gamma\right)-3H+x\frac{\mathcal{C}^{(h)}}{\sqrt{N}},$ (23)
    $\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}h_{i}\sigma_{i}^{a}(t=1)\underset{N\to\infty}{=}\mathrm{erf}\left[\dfrac{\frac{3}{4}(1-2\lambda)+H}{\sqrt{2[\left(\gamma+\frac{9}{16}\right)(1-2\lambda)^{2}]}}\right]\;\;\;\;\mathrm{with}\;\;a=1,\cdots,3.$ (24)
@@
The ranges of stability suggested by these analytical computations are presented in Fig. 2 and corroborated by numerical tests. From this analysis it turns out that the configuration $\bm{\sigma}^{(1,2,3)}$ we are interested in is stable for relatively small values of $\lambda$ and of $H$, corresponding to the region highlighted by the green crosses in Fig. 2 (first row). However, this state represents only a local minimum in the energy landscape and, if we initiate the dynamics from a different initial state, we may no longer converge to $\bm{\sigma}^{(1,2,3)}$, as shown in Fig. 2 (second to fourth rows). Also, in this noiseless scenario, the configuration $\bm{\sigma}^{(h)}$ turns out to be stable for any choice of the parameters $\lambda$ and $H$, thus, some degree of noise is in order for this model to disentangle mixtures. This constitutes an analogy with the standard Hopfield model, where odd mixtures like $\textrm{sign}(\bm{\xi}^{1}+\dots+\bm{\xi}^{2n+1})$, with $n\in\mathbb{N}$, result to be stable at sufficiently low temperatures, thus the application of a certain degree of noise ($\beta^{-1}>0$) is a useful strategy to avoid these “errors” [^7] The beneficial role of thermal noise, when dealing with mixtures in the Hopfield model, was emphasized by Daniel Amit in his seminal work Amit .. Here the configuration $\bm{\sigma}^{(h)}=\textrm{sign}(\bm{\xi}^{1}+\dots+\bm{\xi}^{L})$ is as well a fixed point and the application of some noise allows the system to escape its attractiveness and possibly move towards $\bm{\sigma}^{(1,\dots,L)}$.
@@ 3.2 Stability analysis in the low-load, noisy, and zero-field regime
In this section, we set $\gamma=0$ and $H=0$, and investigate the stability of two possible solutions of the saddle-point equations (11), that is, $\bm{\sigma}^{(h)}$ and $\bm{\sigma}^{(1,2,3)}$, corresponding to, respectively, the input and the target output of the disentanglement task under study. More precisely, we apply the fixed-point iteration technique to (11), by starting the procedure with the configurations $\bm{\sigma}^{(h)}$ and $\bm{\sigma}^{(1,2,3)}$. The related solutions are denoted by $\bar{\bm{m}}^{(h)}\in[-1,+1]^{K\times L}$ and $\bar{\bm{m}}^{(1,2,3)}\in[-1,+1]^{K\times L}$ and depicted in Fig. 3.
@@
are fixed points for the equation (11), with the scalars $m^{\prime}$ and $m^{\prime\prime}$ depending, in general on $\beta$ and $\lambda$. As $\beta^{-1}\to 0$, $m^{\prime}=1$ and $m^{\prime\prime}=0.5$, in such a way that $\bar{\bm{m}}_{\{\mu\leq L\}}^{(1,2,3)}$ and $\bar{\bm{m}}_{\{\mu\leq L\}}^{(h)}$ sharply correspond to $\bm{\sigma}^{(1,2,3)}$ and $\bm{\sigma}^{(h)}$, while, as $\beta^{-1}$ is increased, $m^{\prime}$ and $m^{\prime\prime}$ progressively decrease, yet the matrix structure (scalar and constant) is fairly preserved; then, beyond a certain value of $\beta^{-1}$, we fail to find a solution with that kind of structure. This failure implies that extremal points nearby $\bm{\sigma}^{(1,2,3)}$ or $\bm{\sigma}^{(h)}$ (according to the magnetization matrix used for the initialization) no longer exist. Remarkably, for a given $\lambda$ (e.g., $\lambda=0.2$) and spanning over larger and larger values of $\beta^{-1}$, this singularity occurs first for the input configuration ($\beta^{-1}\approx 0.45$) and then for the output configuration ($\beta^{-1}\approx 0.55$).

### v2

@@ 4 Disentangling spurious states
The modular structure of an $L$-directional associative memory, can be leveraged to tackle several kinds of task beyond standard pattern retrieval. For instance, in AABCR-2024 , we considered pattern disentanglement in the case where patterns retrievable by different layers were independent. Dropping this independence condition makes the task more challenging and, in the current work, we deepen such a scenario. Specifically, we aim to exploit the model (5) for disentangling spurious states, that is, we want to input information in the form of a mixtures of $L$ patterns (without loss of generality we consider the first $L$ patterns) as $\textrm{sign}(\bm{\xi}^{1}+\bm{\xi}^{2}+...+\bm{\xi}^{L})$ and to get as output the single components $\bm{\xi}^{1},\bm{\xi}^{2},...,\bm{\xi}^{L}$, one per layer. In other words, we want the configurations $\bm{\sigma}^{a}=\bm{\xi}^{a}$ for a=1,,La=1,...,L (or any permutation that ensures that different layers retrieve all the different patterns in the input), referred to as $\bm{\sigma}^{(1,2,...,L)}$, to be stable and attracting the configuration $\bm{\sigma}^{a}=\textrm{sgn}(\bm{\xi}^{1}+\bm{\xi}^{2}+...+\bm{\xi}^{L})$ for a=1,,La=1,...,L. Given this task, a natural choice for the field acting on each layer is
$h_{i}=\mathrm{sign}\left(\sum_{\mu=1}^{L}\xi_{i}^{\mu}\right),\penalty 10000\ \penalty 10000\ \textrm{for}\penalty 10000\ \penalty 10000\ i=1,...,N.$ (12)
Notice that this field is layer independent hence the superscript $a$ has been dropped.
The evolution towards the target configuration $\bm{\sigma}^{(1,2,...,L)}$ can be checked by different means. In particular, in Secs. 4.1-4.2, we analytically investigate whether the latter corresponds to the stationary configuration resulting from the self-consistent equations (11), when the fields (12) are applied. Next, in Secs. 4.3-4.4, we numerically investigate whether, starting from the input configuration $\bm{\sigma}^{a}=\bm{h}$ for a=1,,La=1,...,L, referred to as $\bm{\sigma}^{(h)}$, and applying the stochastic local-field-alignment (see, e.g., Amit ), the system eventually reaches the target configuration and this is stable. We recall that the stochastic local-field-alignment plays as neural dynamics for the network and reads
@@
$\tilde{h}_{i}^{a}(t)$ $=$ $\dfrac{1}{N}\displaystyle\sum\limits_{b=1}^{L}g_{ab}\displaystyle\sum\limits_{\mu=1}^{K}\displaystyle\sum\limits_{j=1}^{N}\xi_{j}^{\mu}\sigma_{j}^{b}(t)\xi_{i}^{\mu}+Hh_{i}$ (14)
where $t$ denotes the time step, $\zeta_{i}^{a}(t)$ is a stochastic contribution [^8] We will set $\zeta=\textrm{atanh}(x)$ with $x$ a uniform random variable ranging in [1,+1][-1,+1]; this choice ensures that the dynamics (13) yields to a Boltzmann-Gibbs stationary state, such that this network can be seen as a generalized Boltzmann machine Coolen . and $\bm{\tilde{h}}^{a}\in\mathbb{R}^{N}$ is the local field acting on neurons in the $a$-th layer (stemming from the interactions with other neurons and from the external field).
@@ 4.1 Stability analysis in the high-load, noiseless regime
As mentioned in Sec. 2, the configuration $\bm{\sigma}^{(1,2,...,L)}$ where different layers retrieve different patters is only possible minimum (out of many) for the cost function (5). Thus, before inspecting the ability of the model to disentangle spurious states, it is worth taking a look at some representative extremal configurations and at their stability in the noiseless scenario ($\beta\to\infty$). Keeping $L=3$, we focus on the following classes of neuronal configurations [^9] We are referring to “classes” of neural configurations, because, beyond the degeneracy due to the permutation of the three patterns over the three layers, there is also a degeneracy due to the symmetry of the cost function (5) under spin flip of all the three layers.:
@@
For each of them we will estimate the energy, the consistency and the stability for a fixed value of the network storage $\gamma\in\mathbb{R}^{+}$. Before proceeding, a couple of remarks are in order. First, the previous neuronal states have been chosen because they are recognized to minimize at least one of the contributions making up the cost function (5) and, in fact, we checked that they are also solutions of the self-consistency equations. However, we recall that the last condition only ensures that these configurations are extremal for the free energy, but not necessarily minima, that is, stable points. Second, here the consistency analysis is pursued by recalling the stochastic dynamics (13), setting $\beta^{-1}=0$, and checking whether the configurations remain unchanged, that is, recasting (13) into an evolutionary rule for the Mattis magnetizations
$m^{a}_{\mu}(t+1)=\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\xi_{i}^{\mu}\sigma_{i}^{a}(t+1)=\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\xi_{i}^{\mu}\sigma_{i}^{a}(t)\mathrm{sign}\left[\tilde{h}_{i}^{a}\sigma_{i}^{a}(t)\right]\penalty 10000\ \penalty 10000\ \textrm{for}\penalty 10000\ \penalty 10000\ a=1,...,L,$ (15)
@@
The analytical estimates for the one-step magnetization (15) and for the energy (5) related to the four configurations above follow from straightforward but pretty lengthy calculations that are detailed in App. C, while here we provide a summary of the results in Figure 2.
@@
Figure 2: We initialize the system in the configuration $\bm{\sigma}^{(1,2,3)}$ (first line), $\bm{\sigma}^{(1,1,1)}$ (second line), $\bm{\sigma}^{(1,1,1^{\prime})}$ (third line), and $\bm{\sigma}^{(h)}$ (fourth line) and we evaluate analytically the related one-step magnetization (15), thereby deriving the stability region (black line) in the $(H,\lambda)$ plane for that solution. Specifically, these lines are obtained by setting $\gamma=0.01$ and by determining in which region of the plane the one-step magnetization (i.e., the error functions in, respectively, eqs. (C.2), (C.4), (C.6), (C.7), (C.9)), exceed a certain threshold, which we set to $0.95$; for $\bm{\sigma}^{(h)}$ no boundaries are detected in the region under consideration. The shade in the color accounts for the energy associated to the related fixed point: the smaller the energy and the darker the color, see also eqs. (C.1), (C.3), (C.5), (C.8). Thus, for small $H$, although $\bm{\sigma}^{(1,2,3)}$ turns out to be stable, its energy is relatively close to zero. These analytical predictions are validated against computational results in order to assess the configuration stabilities versus small perturbations. To this purpose, we initialize the system in a configuration obtained from $\bm{\sigma}^{(1,2,3)}$ (first line), from $\bm{\sigma}^{(1,1,1)}$ (second line), from $\bm{\sigma}^{(1,1,1^{\prime})}$ (third line), and from $\bm{\sigma}^{(h)}$ (fourth line), by flipping randomly its entries: the flip is implemented by multiplying each neuron variable $\sigma_{i}^{a}$ by a random variable $\chi_{i}^{a}$ drawn from $P(\chi)=\frac{1+r}{2}\delta(\chi-1)+\frac{1-r}{2}\delta(\chi+1)$, where $r=1.0$ (left column), $r=0.8$ (middle column), and $r=0.5$ (right column), clearly, the larger $r$ and the closer the initial configuration to the reference. Then, we implement the dynamics (13) with $T=0$, up to convergence to a fixed point. This is repeated for several choices of the parameters $H$ and $\lambda$ sampled uniformly in, respectively, $[0,2]$ and $[0,0.5]$ and for fixed $N=5000$ and $K=50$. Different final states are recorded and represented by different symbols and colors, as reported by the legend on the right: $\bm{\sigma}^{(1,2,3)}$ (green $\times$), $\bm{\sigma}^{(1,1,1)}$ (blue ++), $\bm{\sigma}^{(1,1,1^{\prime})}$ (magenta $\circ$), $\bm{\sigma}^{(h)}$ (red $\triangle$), or none of those considered in this section (gray $\bullet$). The patterns presented in the figure are just for illustrative purposes as both analytical and numerical results are obtained for a Rademacher dataset; for an analysis involving structured data we refer to Sec. $4.$
@@
From this analysis it turns out that the configuration $\bm{\sigma}^{(1,2,3)}$ we are interested in is stable for relatively small values of $\lambda$ and of $H$, corresponding to the region highlighted by the green crosses in Figure 2 (first row). However, this state represents only a local minimum in the energy landscape and, if we initiate the dynamics from a different initial state, we may no longer converge to $\bm{\sigma}^{(1,2,3)}$, as shown in Figure 2 (second to fourth rows). Also, in this noiseless scenario, the configuration $\bm{\sigma}^{(h)}$ turns out to be stable for any choice of the parameters $\lambda$ and $H$, thus, some degree of noise is in order for this model to disentangle mixtures. This constitutes an analogy with the standard Hopfield modules, where odd mixtures like $\textrm{sign}(\bm{\xi}^{1}+\dots+\bm{\xi}^{2n+1})$, with $n\in\mathbb{N}$, result to be stable at sufficiently low temperatures, thus the application of a certain degree of noise ($\beta^{-1}>0$) is a useful strategy to avoid these “errors” Amit . Here the configuration $\bm{\sigma}^{(h)}=\textrm{sign}(\bm{\xi}^{1}+\dots+\bm{\xi}^{L})$ is as well a fixed point and the application of some noise allows the system to escape its attractiveness and possibly move towards $\bm{\sigma}^{(1,\dots,L)}$. The previous analysis was carried out for a fixed value of the network storage $\gamma=0.01$, however, extensive simulations were performed for different values of relatively low values of $\gamma$, demonstrating the overall robustness of the network’s dsentangling capability with respect to variations in network storage (see also App. G).
@@ 4.2 Stability analysis in the low-load, noisy, and zero-field regime
In this section, we set $\gamma=0$ and $H=0$, and we focus on two possible solutions of the saddle-point equations (11), that is, $\bm{\sigma}^{(h)}$ and $\bm{\sigma}^{(1,2,3)}$, corresponding to, respectively, the input and the target output of the disentanglement task under study. More precisely, we apply the fixed-point iteration technique to (11), by starting the procedure with the configurations $\bm{\sigma}^{(h)}$ and $\bm{\sigma}^{(1,2,3)}$. The related solutions are denoted by $\bar{\bm{m}}^{(h)}\in[-1,+1]^{K\times L}$ and $\bar{\bm{m}}^{(1,2,3)}\in[-1,+1]^{K\times L}$ and depicted in Figure 3.
@@
are fixed points for the equation (11), with the scalars $m^{\prime}$ and $m^{\prime\prime}$ depending, in general, on $\beta$ and $\lambda$. As $\beta^{-1}\to 0$, $m^{\prime}=1$ and $m^{\prime\prime}=0.5$, in such a way that $\bar{\bm{m}}_{\{\mu\leq L\}}^{(1,2,3)}$ and $\bar{\bm{m}}_{\{\mu\leq L\}}^{(h)}$ sharply correspond to the magnetizations related to the configurations $\bm{\sigma}^{(1,2,3)}$ and $\bm{\sigma}^{(h)}$, while, as $\beta^{-1}$ is increased, $m^{\prime}$ and $m^{\prime\prime}$ progressively decrease, yet the matrix structure (scalar and constant) is fairly preserved; then, beyond a certain value of $\beta^{-1}$, we fail to find a solution with that kind of structure. This failure implies that extremal points nearby $\bm{\sigma}^{(1,2,3)}$ or $\bm{\sigma}^{(h)}$ (according to the magnetization matrix used for the initialization) no longer exist. Remarkably, for a given $\lambda$ (e.g., $\lambda=0.2$) and spanning over larger and larger values of $\beta^{-1}$, this singularity occurs first for the input configuration $\bm{\sigma}^{(h)}$ ($\beta^{-1}\approx 0.45$) and then for the output configuration $\bm{\sigma}^{(1,2,3)}$ ($\beta^{-1}\approx 0.55$).